Theodore needs to mix a 20% acid solution with a 60% acid solution to create 200 millileters of a 24% solution. How many millileters of each solution must Theodore use?
Answer by stanbon(75887)

Theodore needs to mix a 20% acid solution with a 60% acid solution to create 200 millileters of a 24% solution. How many millileters of each solution must Theodore use?
-------------------
Equation:
acid + acid = acid
0.20x + 0.60(200-x) = 0.24*200
-----
20x + 60*200 - 60x = 24*200
-40x = -36*200
----
x = 5*36
x = 180 ml (amt. of 20% solution needed)
200-x = 20 ml (amt. of 60% solution needed)
--------------------
Cheers,
Stan H.
----------------